find 3 consecutive intergers whose sum is -33
Answer by jim_thompson5910(35256)


x+(x+1)+(x+2) = -33

3x+3 = -33

3x = -33-3

3x = -36

x = -36/3

x = -12

x+1 = -11

x+2 = -10

So the three consecutive integers are -12, -11, -10